OpenStudy (anonymous):

of the 15 students who solved the challenge question in mathematics contest. 8 were enrolled in mathematics of data mangement. Five of the solutions are selected at random for a display a) Prepare a table of the probablility distribution for thenumber of solutions in the display that were prepared by mathematics of data students.

OpenStudy (anonymous):

the possible number of students selected from the 8 enrolled in data management are x = 0 x = 1 x= 2 x = 3 x = 4 x = 5 the number of ways to select none is \[P(x=0)=\frac{7\times 6\times 5\times 4\times 3}{\dbinom{15}{5}}\]

OpenStudy (anonymous):

\[P(x=1)=\frac{\dbinom{8}{1}\times \dbinom{7}{4}}{\dbinom{15}{5}}\] that is the number of ways to choose one from the 8 times the number of ways to choose 4 from the other seven, divided by total number of ways to choose 5 from 15 others are similar

OpenStudy (anonymous):

Why is it only up to x=5? Why not x=8? since there are 8 people?